3,5,8,13,22,? which no. at ?

3,5,8,13,22,?
which no. at ?

Grade:11

3 Answers

Sher Mohammad IIT Delhi
10 years ago
number added to get next element 2 3 5 8 13
35 will be next

sher mohmmad
B.Tech, IIT Delhi
srinu vasa rao
10 years ago
35 it`s a febbinoci series....................... srinu iiiit rkv
Anand Ramesh
10 years ago
39 3+(1+2^0)=5, 5+(1+2^1)=8, 8+(1+2^2)=13, 13+(1+2^3)=22, 22+(1+2^4)=39... Anand Ramesh XII INDIRA GANDHI PUBLIC SCHOOL MAMBARAM